5 g ice at 0°C is mixed with 1 g steam at 100° C. Find the final temperature and composition of the mixture.

Grade:11

1 Answers

Arun
5 years ago
steam and ice will first convert into liquid water at 100C and 0C respectively. Then they will form a mixture whose resultant temperature has been asked. 

heat lost by steam = heat gained by ice 
[1 * 2268] + [1 * 4.2 * (100 - t)] = [5 * 336] + [5 * 4.2 * (t - 0)] 

{ here t is the final temperature of the mixture, 336 is the specific latent heat of melting for ice and 2268 is the specific latent heat of evaporation for steam } 
 
[ 2268 + 420 - 4.2t ] = 5[ 336 + 4.2t ]  
=> 2688 - 4.2t = 1260 + 21 t 
=> 1428 = 21 t + 4.2t 
=> 1428 = 25.2 t
=> 1428/25.2 = t 
=> 56.66= t